Grade 12th passMechanics

The coefficient of static friction between Teflon and scrambled bag eggs is
about 0.04. What is the smallest angle from the horizontal that will cause the eggs to slide
across the bottom of a Teflon-coated skillet?

Uzma
10 Years agoGrade 12th pass

1 Answer

Khaled Hasan
8 Years ago
Solution: Free-body diagram when egg just starts to slip

Resolve into components:
 
perpendicular to pan     W⊥ = W cos θ
 parallel to pan            W║ = W sin θ
 
When the egg is just about to slip, there is no net force in either direction, so
 perpendicular to pan     N – W⊥ = 0..... (1)
 parallel to pan               f – W║ = 0...... (2)
 
Equation (1)
⇒     N – W cos θ = 0
 so    cos θ = N / W
 
Equation (2) ⇒      
μsN – W sin θ = 0
so         sin θ = μsN / W
 
Now        tan θ = sin θ / cos θ = μsN / W  ×  W / N 
                                            = μs = 0.04
so           θ = tan–1(0.04) ! ! !   = 2.3°